I am trying to find the measure of an angle whose supplement is 4 times the measure of its complement. I am so confused. Please help me!
Answer by uma(370)

Let the angle be x
Its supplement = 180 - x
Its complement = 90 - x
As per the given data,
4(90 - x) = 180 - x
==> 360 - 4x = 180 - x
==> 360 - 180 - 4x = -x
==> 180 - 4x = - x
==> 180 = 4x - x
==> 180 = 3x
==> 180/3 = x
==> x = 60

Thus the angle is 60degrees whose supplement is 4 times the measure of its complement.
Good Luck!!!